find equation of the line through points (-5,3) snf (6,1) test question
Answer by checkley77(12844)

(-5,3) (6,1)
slope=(1-3)/(6+5)
slope=-2/11
3=-5(-2/11)+b
3=10/11+b
b=3-10/11
b=(33-10)/11
b=23/11 answer for the y intercept.
the formula is:
y=-2x/11+23/11